OpenStudy (karinamtz0213):

Given a parabola with zeros at x=0 and x=4, which of the following MUST be true: I. The parabola opens upward II. The parabola has a line of symmetry at x=2 III. The parabola intercepts the y-axis at y=0 a. I only b. III only c. I and III only d. II and III only e. I, II, and III

OpenStudy (whpalmer4):

Does this parabola open upward or downward? \(y = x^2\) How about \(y = -x^2\)?

OpenStudy (karinamtz0213):

it opens upward

OpenStudy (whpalmer4):

yes, \(y = x^2\) opens upward. but what about \(y = -x^2\)?

OpenStudy (karinamtz0213):

that i do not know

OpenStudy (whpalmer4):

Note that this is \(y = -1*x^2\) not to be confused with \(y = (-x)^2\)

OpenStudy (whpalmer4):

Well get out a piece of paper and plot three points: x = -1, x = 0, x = 1 (using \(y = -x^2\) to find the corresponding \(y\) values

OpenStudy (karinamtz0213):

then?

OpenStudy (whpalmer4):

then you should be able to tell me if the parabola \(y = -x^2\) opens upward or downward.

OpenStudy (whpalmer4):

Well?